sipho and litha are brothers . the sum of their ages is 29 . in 6 year's time sipho will be the same age as litha was 3 years ago

1
Expert's answer
2022-03-01T14:48:24-0500

Let Sipho and Litha's present ages be x years and y years respectively.


We have that


"x+y=29" -----------> (1)


"x+6=y-3"


"=>x-y=-9" -----------> (2)


Adding equation (1) and (2) together, we have;


"2x=20"


"=> x = 10"

Put x=10 in equation (2)


"10-y=3"


"=> y=10-3=7"


Therefore, Sipho and Litha's present ages are 10years and 7years respectively.